Upstream, downstream. Junior’s boat will go 15 miles per hour in still water. If he can go 12 miles downstream in the same amount of time as it takes to go 9 miles upstream, then what is the speed of the current?
Answer by stanbon(75887)

Junior’s boat will go 15 miles per hour in still water. If he can go 12 miles downstream in the same amount of time as it takes to go 9 miles upstream, then what is the speed of the current?
------------
Downstream DATA;
distance = 12 miles; rate = 15+c mph ; time = d/r = 12/(15+c) hrs.
-----------------
Upstream DATA:
distance = 9 miles ; rate = 15-c mph ; time = d/r = 9/(15-c) hrs.
---------------------------------------
Time = time
12/(15+c) = 9/(15-c)
12(15-c) = 9(15+c)
12*15 - 12c = 9*15 + 9c
21c = 3*15
c = 3*15/21
c = 45/21 = 2.1429..(speed of the current)
Cheers,
Stan H.